# A Unified Spectral Framework for Aging, Heterogeneous, and Distributed Order Systems via Weighted Weyl-Sonine Operators

### v1

@@ 4.2 Application: Weighted Distributed Order and Ultra-Slow Transitions
A significant advantage of the Weyl-Sonine framework is its ability to encapsulate operators of distributed order. These operators model complex multi-scale memory effects and "ultra-slow" diffusion processes where the mean squared displacement grows logarithmically rather than as a power law.
@@
Let $b:[0,1]\to[0,\infty)$ be a probability density function representing the distribution of differentiation orders. We define the Weighted Distributed Order Operator, $\mathbb{D}_{\psi,\omega}^{b}$, as the integral superposition of the weighted fractional derivatives:
@@
It is straightforward to show that $\mathbb{D}_{\psi,\omega}^{b}$ belongs to the Sonine class defined in Section 3. Specifically, it corresponds to the effective kernel $\kappa_{\text{eff}}(t)=\int_{0}^{1}\frac{b(\alpha)}{\Gamma(1-\alpha)}t^{-\alpha}\,d\alpha$.
Crucially, since $s^{\alpha}$ is a Bernstein function for $\alpha\in(0,1)$ and $b(\alpha)\geq 0$, the effective spectral symbol
$\Phi_{b}(s)=\int_{0}^{1}b(\alpha)s^{\alpha}\,d\alpha$
is itself a Complete Bernstein Function. This guarantees that the distributed order operator preserves the positivity and monotonicity properties established in our main theorems.
@@ 4.2.1 Example: Retarded Aging via Logarithmic Diffusion
Consider the canonical case of a uniform distribution of orders, $b(\alpha)=1$. The spectral symbol is computed directly:
$\Phi_{\text{uni}}(i\xi)=\int_{0}^{1}(i\xi)^{\alpha}\,d\alpha=\frac{i\xi-1}{\ln(i\xi)}.$ (37)
Applying Theorem 3.14, the fundamental solution (Green’s function) for the associated relaxation problem is governed by the multiplier:
$\widehat{\mathcal{G}}(\xi)=\frac{1}{\Phi_{\text{uni}}(i\xi)+\lambda}=\frac{\ln(i\xi)}{i\xi-1+\lambda\ln(i\xi)}.$ (38)
@@ 5 Conclusions and Perspectives
In this work, we have successfully constructed a unified mathematical framework that bridges the gap between General Fractional Calculus and the analysis of heterogeneous, aging media. By defining the Weighted Weyl-Sonine operators, we have extended the scope of non-local theory to the entire real line $\mathbb{R}$, allowing for memory kernels of arbitrary nature—governed by Complete Bernstein Functions—while simultaneously incorporating time-dependent material properties via the admissible structure pair $(\psi,\omega)$.
@@
The versatility of this approach was illustrated through the solvability of generalized relaxation models. We derived explicit Green’s functions for arbitrary admissible symbols, ensuring the preservation of positivity in the resulting dynamics. In particular, we demonstrated that our framework naturally encompasses Weighted Distributed Order Operators, proving that phenomena of ultra-slow diffusion and retarded aging can be rigorously analyzed and solved using the distributed spectral symbol $\Phi_{b}(\xi)$.
@@ B.2 Definition via Parseval’s Identity
This definition is consistent with the Plancherel isometry on $L^{2}_{\psi,\omega}$. It ensures that for any convolution kernel $k$ (viewed as a distribution), the spectral mapping property holds rigorously:
$\mathcal{F}_{\psi,\omega}\left[\mathfrak{D}_{\psi,\omega}^{(k)}u\right]=\mathcal{F}_{\psi,\omega}[k]\cdot\mathcal{F}_{\psi,\omega}[u]$
in the sense of distributions, where the product is understood as the multiplication of a distribution by a smooth function (the spectral symbol) or via the exchange formula.

### v2

@@ 4.2 Application: Weighted Distributed Order and Ultra-Slow Transitions
A significant advantage of the Weyl-Sonine framework is its ability to encapsulate operators of distributed order. These operators model complex multi-scale memory effects and "ultra-slow" diffusion processes where the mean squared displacement grows logarithmically rather than as a power law (a behavior distinct from the smooth crossover to normal diffusion depicted in Figure 2).
Figure 2: Diffusive Regimes. Mean Squared Displacement (MSD) for different memory structures. The plot illustrates the transition from subdiffusion ($\alpha<1$) to normal diffusion ($\alpha=1$) typical of tempered systems. Note that the retarded aging process discussed in the text follows an ultra-slow logarithmic trajectory (flatter than the curves shown here).
@@
Let $b:[0,1]\to[0,\infty)$ be a probability density function representing the distribution of differentiation orders. We define the Weighted Distributed Order Operator, $\mathbb{D}_{\psi,\omega}^{b}$, as the integral superposition:
@@
This operator belongs to the Sonine class with order $n=1$. Its associated memory kernel is $\kappa_{\text{eff}}(t)=\int_{0}^{1}\frac{b(\alpha)}{\Gamma(1-\alpha)}t^{-\alpha}\,d\alpha$. To maintain consistency with Theorem 4.2, we define the Spectral Symbol $\Phi_{b}(s)$ as the transform of this kernel:
$\Phi_{b}(s)=\mathcal{L}\{\kappa_{\text{eff}}\}(s)=\int_{0}^{1}b(\alpha)s^{\alpha-1}\,d\alpha.$ (37)
Note that while the intuitive symbol of the operator is $\int b(\alpha)s^{\alpha}d\alpha$, our structural parameter $\Phi_{b}(s)$ includes the singular factor $s^{-1}$ inherent to the memory kernel definition. Since $s^{\alpha-1}$ is the Stieltjes transform of a measure for $\alpha\in(0,1)$, $\Phi_{b}(s)$ preserves the necessary analytic properties.
@@ 4.2.1 Example: Retarded Aging via Logarithmic Diffusion
Consider the canonical case of a uniform distribution of orders, $b(\alpha)=1$. We compute the spectral symbol according to (37):
$\Phi_{\text{uni}}(i\xi)=\int_{0}^{1}(i\xi)^{\alpha-1}\,d\alpha=\frac{1}{i\xi}\int_{0}^{1}(i\xi)^{\alpha}\,d\alpha=\frac{i\xi-1}{i\xi\ln(i\xi)}.$ (38)
Now we apply Theorem 4.2. The characteristic multiplier for the Green’s function is determined by substituting $\Phi_{\text{uni}}$ into Eq. (35):
$i\xi\Phi_{\text{uni}}(i\xi)+\lambda=i\xi\left(\frac{i\xi-1}{i\xi\ln(i\xi)}\right)+\lambda=\frac{i\xi-1}{\ln(i\xi)}+\lambda.$
Thus, the Green’s function is determined by the inverse Fourier transform of:
$\widehat{\mathcal{G}_{\lambda}}(\xi)=\frac{\ln(i\xi)}{i\xi-1+\lambda\ln(i\xi)}.$ (39)
@@
4.3 Application: Inertial Dynamics and Wave Propagation in Expanding Media
Finally, we address the Oscillatory Regime identified in Remark 3.2. Unlike the relaxation and diffusion processes discussed above, inertial dynamics (such as wave propagation) require memory kernels that allow for energy exchange rather than pure dissipation.
Consider the generalized telegrapher’s equation in a time-varying medium. In our framework, this is modeled by selecting the Bessel kernel from Table 1 (entry II):
$\kappa_{\text{wave}}(t)=J_{0}(2\sqrt{\gamma t}),\quad\gamma>0.$ (40)
This kernel is not a Bernstein function; it is signed and oscillatory. Its spectral symbol is given by the essential singularity:
$\Phi_{\text{wave}}(s)=\mathcal{L}\{J_{0}(2\sqrt{\gamma t})\}(s)=\frac{1}{s}e^{-\gamma/s}.$ (41)
Spectral Analysis of Oscillations: Using the spectral resolution provided by Theorem 4.2, the dynamic behavior is governed by the poles of the Green’s function $\mathcal{G}_{\lambda}$. The characteristic equation (setting the denominator of Eq. (35) to zero) becomes:
$i\xi\Phi_{\text{wave}}(i\xi)+\lambda=i\xi\left(\frac{1}{i\xi}e^{-\gamma/i\xi}\right)+\lambda=e^{i(\gamma/\xi)}+\lambda=0.$ (42)
The dispersion relation $e^{i(\gamma/\xi)}+\lambda=0$ implies that for modes on the unit spectral circle ($|\lambda|=1$), the wavenumbers $\xi$ are real. This creates a spectrum of non-decaying propagating modes (neutral stability), a hallmark of wave mechanics that contrasts sharply with the dissipative spectrum of parabolic operators.
Physical Interpretation of the Weights: The solution $u(t)$ obtained via (34) describes a wave packet. The structural functions play a geometric role:
  • The scale $\psi(t)$ acts as a local clock. If $\psi(t)=t^{\beta}$ with $\beta>1$ (expanding medium), the perceived frequency of the wave undergoes a redshift over time.
  • The weight $\omega(t)$ modulates the amplitude. An exponential weight $\omega(t)=e^{\mu t}$ introduces a generalized friction or amplification term without altering the causality of the wave equation.
Figure 3: Wave Propagation in Expanding Media. (A) Classical inertial dynamics (Telegrapher’s equation) showing constant oscillation frequency. (B) Weyl-Sonine evolution with a time-scale deformation $\psi(t)\sim t^{0.6}$. The wave packet exhibits a cosmological redshift, where the wavelength stretches as the internal clock slows down relative to laboratory time.
This example confirms that the Weighted Weyl-Sonine framework unifies parabolic (diffusive) and hyperbolic (wave-like) phenomena under a single algebraic structure, distinguishing them solely by the spectral class of the memory kernel $\Phi$.
@@ 5 Conclusions and Perspectives
In this work, we have successfully constructed a unified mathematical framework that bridges the gap between General Fractional Calculus and the analysis of heterogeneous, aging media. By defining the Weighted Weyl-Sonine operators, we have extended the scope of non-local theory to the entire real line $\mathbb{R}$, allowing for memory kernels of arbitrary spectral nature—encompassing both the diffusive regime (governed by Complete Bernstein Functions) and the inertial regime (oscillatory kernels)—while simultaneously incorporating time-dependent material properties via the admissible structure pair $(\psi,\omega)$.
@@
The versatility of this approach was illustrated through the solvability of evolution equations across distinct physical landscapes. We derived explicit Green’s functions for arbitrary admissible symbols, distinguishing between monotonic relaxation (where positivity is strictly preserved) and wave propagation (where oscillations arise from singular spectra). Additionally, we demonstrated that our framework naturally encompasses Weighted Distributed Order Operators, proving that phenomena of ultra-slow diffusion and retarded aging can be rigorously analyzed using the distributed spectral symbol $\Phi_{b}(\xi)$.
@@ B.2 Definition via Parseval’s Identity
This definition is consistent with the Plancherel isometry on $L^{2}_{\psi,\omega}$. It ensures that for any convolution kernel $\kappa$ (viewed as a distribution), the spectral mapping property holds rigorously:
$\mathcal{F}_{\psi,\omega}\left[\mathfrak{D}_{\psi,\omega}^{(\kappa)}u\right]=\mathcal{F}_{\psi,\omega}[\kappa]\cdot\mathcal{F}_{\psi,\omega}[u]$
in the sense of distributions, where the product is understood as the multiplication of a distribution by the corresponding spectral symbol (multiplier function) or via the exchange formula.
Declaration of competing interest
The author declares that he has no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of Generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the author(s) used Gemini (Google) in order to improve the readability and language quality of the manuscript, as well as to assist in structuring the discussion of the physical interpretations. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.